# Euclid preparation

### v1

@@ 6.2 Validation of the spectroscopic dispersion law
RGS180 | 0.14 | 0.15 | 0.33 | 0.38 |
Additionally, we extracted the -order position from a template fit, replacing the Fabry–Pérot template with the Argon SED template whenever the -order was visible on the 2D spectrogram. This was primarily possible for acquisitions with the BGS000 grism. Contrary to this, the signal was below the detectable signal-to-noise ratio for the red grisms. Since we did not move the telescope simulator when changing the light source, we assumed for those cases that the -order position of the Argon spectrogram was the same as that of the Fabry–Pérot spectrogram from the previous exposure.
@@
Using calibration parameters that were derived from the Fabry–Pérot spectrogram analysis (see Sect. 6.1), we estimated the position of the Argon emission lines relative to the -order position and compared the predicted positions with the measured position.
The bias resulting from this comparison is reported in Table 1, defined as the average of the distribution as shown in Fig. 13. It is induced by the persistence present in the images that tend to either bias the centroiding of the Argon lines themselves or the estimate of the position of the -order. As was described above, for the BGS000 an Argon SED template was used to evaluate the -order position, while for both red grisms this was taken from the previous exposure, assuming the telescope did not move between exposures. As a consequence, the weakness of the -order and the persistence signal present in the BGS000 spectrogram leads to a larger bias than what is observed for the corresponding distribution of the red grisms. Nonetheless, despite this bias reported in the Table 1, we found that the predicted position of the Argon lines fall within a distance of pixels of the measured lines, which validates our calibration parametrisation.
@@ 7 NISP Spectroscopic performance
For clarity and simplicity, the forthcoming discussion will be confined to the description of the dispersion law averaged over the NISP FoV. Table 2 reports the dispersion coefficients in both the Chebyshev and regular basis and reveals that when the grisms are centred onto their nominal position, the cross-dispersion coefficient is about two orders of magnitude lower than the dispersion coefficient . This aligns with the intended design of the grisms which primarily focuses on dispersing light in a single direction. Optical and mechanical distortions as well as the manufacturing limitations contribute to the distortion of the spectral track in both directions. However, even with those limitations, the higher-order coefficients ( with ) are small enough for the dispersion to be nearly a linear dispersion law mostly directed towards the axis.
@@
In the context of slit-less spectroscopy, spectral resolution is defined by a combination of both the intrinsic instrument PSF as well as the characteristics of the targeted object’s size, shape, and surface brightness distribution. Referring to the definition provided in Euclid’s scientific requirement, we define NISP’s resolving power as
In the above equation, corresponds to the per-pixel spectral dispersion along the dispersion direction444along the -axis for RGS000 and RGS180 in nominal position and in a mixture of and axis when those grisms are at inclination and is expressed in /px. The corresponds to the full-width-half-maximum of the target source with an axisymmetric Gaussian profile of effectives variance . This effective full-width-half-maximum results from the convolution between the luminosity profile of the source and the PSF and can be expressed as:
The is the full-width-half-maximum of the PSF, in pixels, at a given wavelength , and corresponds to the full-width-half-maximum, in pixels, of the target and depends on the source’s surface brightness distribution.
@@
Figure 14 presents the variation across the field-of-view of the dispersion coefficients at the mid-bandpass wavelength, for all combinations of grism/rotation positions that were tested. The dispersion smoothly varies across the NISP FoV following a radial gradient with its maximum, or minimum, on the left-hand side of the focal plane at about mosaic coordinates . Notice the difference between the two red grisms: RGS000 has an average dispersion of the order of while RGS180’s average dispersion is of the order of . The negative sign of the RGS180 dispersion originates from its design which imposes a dispersion of the light in the opposite direction from RGS000. But on overall, RGS000 and RGS180 have a very similar absolute dispersion in term of value and profile over the focal plane array, leading to a similar performance for both grisms. Note that, to highlight these similarities and due to the negative dispersion value of RGS180, color scale for RGS180 is reversed relative to that of RGS000. In contrast, BGS000 has a lower dispersion with an average value of the order of . All these values are compatible with the design-specifications for each grism.
Combining this estimate of the dispersion with the measured PSF size, NISP’s resolving power can be derived using Eqs. 6 and 7. The result is summarized in Fig. 15 which shows NISP’s averaged resolving power for a source of 05 FWHM. The averaged resolving power from the blue grism increases from about at to at , while both red grisms see their resolving power increase from at to at . These values have to be compared to the NISP scientific requirements of a resolving power for the blue grism and for the red grisms and for a typical source size of . The combination of the high optical quality of the grisms and the NISP camera and collimator lenses ensures that NISP fulfils its requirements in term of spectral dispersion and resolving power.
Again, shorter test were conducted at the payload module level with the addition of the Euclid mirrors and dichroic element to the light path. This allowed us to verify that the addition of the folding mirrors as well as the transmission of the near-infrared light beam through the dichroic did not significantly modify NISP’s spectral dispersion and resolving power.
@@ 8 Conclusions and available data products
While a non-conformity was revealed for the grism RGS270, which resulted in the modification of the Euclid survey strategy, the analysis of the ground dataset presented in this work shows that NISP’s grisms BGS000, RGS000, and RGS180 have an extremely high image quality with a PSF encircled energy size well below the scientific requirements and very close to that of an ideal, solely diffraction-limited instrument.
Secondly, these tests allowed us to verify that the spectral dispersion of NISP lies above requirements and showed that NISP’s resolving power is larger than from to . The RGS000, RGS180, RGS270, and BGS000 throughput for the and spectral orders at the centre of the NISP focal plane array are available at an ESA server555https://euclid.esac.esa.int/msp/refdata/nisp/NISP-SPECTRO-PASSBANDS-V1. Available throughput numbers include contributions from the grisms, the telescope, and NI-OA, as well as the mean detector quantum efficiency. For convenience, the tables are available in both ASCII and FITS format.
@@ Appendix A On ground calibration coefficients
Table 3 to 9 give the calibration coefficients for Eq. 3, used to model special dependency of the coefficient needed in Eq. 2.
Grism | Coef. | Chebychev polynomial coefficients | Regular polynomial coefficient | ||||||

### v2

@@ 5.2 Validation of the spectroscopic dispersion law
Additionally, we extracted the -order position from a template fit, replacing the Fabry–Pérot template with the Argon spectral energy distribution (SED) template whenever the -order was visible on the 2D spectrogram. This was primarily possible for acquisitions with the BGS000 grism. Contrary to this, the signal was below the detectable signal-to-noise ratio for the red grisms. Since we did not move the telescope simulator when changing the light source, we assumed for those cases that the -order position of the Argon spectrogram was the same as that of the Fabry–Pérot spectrogram from the previous exposure.
@@
| RGS180 | 0.14 | 0.15 | 0.33 | 0.38 | ||
@@
Using calibration parameters that were derived from the Fabry–Pérot spectrogram analysis (see Sect. 5.1), we estimated the position of the Argon emission lines relative to the -order position and compared the predicted positions with the measured position.
The bias resulting from this comparison is reported in Table 1, defined as the average of the distribution as shown in Fig. 13. It is induced by the persistence present in the images that tend to either bias the centroiding of the Argon lines themselves or the estimate of the position of the -order. As was described above, for the BGS000 an Argon spectral energy distribution template was used to evaluate the -order position, while for both red grisms this was taken from the previous exposure, assuming the telescope did not move between exposures. As a consequence, the weakness of the -order and the persistence signal present in the BGS000 spectrogram leads to a larger bias than what is observed for the corresponding distribution of the red grisms. Nonetheless, despite this bias reported in the Table 1, we found that the predicted position of the Argon lines fall within a distance of pixels of the measured lines, which validates our calibration parametrisation.
@@ 6 NISP Spectroscopic performance
For clarity and simplicity, the forthcoming discussion will be confined to the description of the dispersion law averaged over the NISP field-of-view. Table 2 reports the dispersion coefficients in both the Chebyshev and regular basis and reveals that when the grisms are centred onto their nominal position, the cross-dispersion coefficient is about two orders of magnitude lower than the dispersion coefficient . This aligns with the intended design of the grisms which primarily focuses on dispersing light in a single direction. Optical and mechanical distortions as well as the manufacturing limitations contribute to the distortion of the spectral track in both directions. However, even with those limitations, the higher-order coefficients ( with ) are small enough for the dispersion to be nearly a linear dispersion law mostly directed towards the axis.
@@
In the context of slitless spectroscopy, spectral resolution is defined by a combination of both the intrinsic instrument point spread function as well as the characteristics of the targeted object’s size, shape, and surface brightness distribution. Referring to the definition provided in Euclid’s scientific requirement, we define NISP’s resolving power as
In the above equation, corresponds to the per-pixel spectral dispersion along the dispersion direction444along the -axis for RGS000 and RGS180 in nominal position and in a mixture of and axis when those grisms are at inclination and is expressed in /px. The corresponds to the full-width-half-maximum of the target source with an axisymmetric Gaussian profile of effectives variance . This effective full-width-half-maximum results from the convolution between the luminosity profile of the source and the point spread function and can be expressed as:
The is the full-width-half-maximum of the point spread function, in pixels, at a given wavelength , and corresponds to the full-width-half-maximum, in pixels, of the target and depends on the source’s surface brightness distribution.
@@
Figure 14 presents the variation across the field-of-view of the dispersion coefficients at the mid-bandpass wavelength, for all combinations of grism/rotation positions that were tested. The dispersion smoothly varies across the NISP field-of-view following a radial gradient with its maximum, or minimum, on the left-hand side of the focal plane at about mosaic coordinates . Notice the difference between the two red grisms: RGS000 has an average dispersion of the order of while RGS180’s average dispersion is of the order of . The negative sign of the RGS180 dispersion originates from its design which imposes a dispersion of the light in the opposite direction from RGS000. But on overall, RGS000 and RGS180 have a very similar absolute dispersion in terms of value and profile over the focal plane array, leading to a similar performance for both grisms. Note that, to highlight these similarities and due to the negative dispersion value of RGS180, colour scale for RGS180 is reversed relative to that of RGS000. In contrast, BGS000 has a lower dispersion with an average value of the order of . All these values are compatible with the design-specifications for each grism.
Combining this estimate of the dispersion with the measured point spread function size, NISP’s resolving power can be derived using Eqs. 6 and 7. The result is summarized in Fig. 15 which shows NISP’s averaged resolving power for a source of 05 Full-Width-Half-Maximum. The averaged resolving power from the blue grism increases from about at to at , while both red grisms see their resolving power increase from at to at . These values have to be compared to the NISP scientific requirements of a resolving power for the blue grism and for the red grisms and for a typical source size of . The combination of the high optical quality of the grisms and the NISP camera and collimator lenses ensures that NISP fulfils its requirements in terms of spectral dispersion and resolving power.
In addition to the measurements described in this work, ground tests were conducted to verify compliance with stray light requirements from in-field and out-of-field sources. These confirmed that parasitic signals are below scientific performance thresholds. No further characterization was done, as the objective was limited to performance verification and time dedicated to the test campaign was constrained.
Again, shorter tests were conducted at the payload module level with the addition of the Euclid mirrors and dichroic element to the light path. This allowed us to verify that the addition of the folding mirrors as well as the transmission of the near-infrared light beam through the dichroic did not significantly modify NISP’s spectral dispersion, resolving power as well as optical performances.
@@ 7 Conclusions and available data products
While a non-conformity was revealed for the grism RGS270, which resulted in the modification of the Euclid survey strategy, the analysis of the ground dataset presented in this work shows that NISP’s grisms BGS000, RGS000, and RGS180 have an extremely high image quality with a point spread function encircled energy size well below the scientific requirements and very close to that of an ideal, solely diffraction-limited instrument.
Secondly, these tests allowed us to verify that the spectral dispersion of NISP lies above requirements and showed that NISP’s resolving power is larger than from to . The RGS000, RGS180, RGS270, and BGS000 throughput for the and spectral orders at the centre of the NISP focal plane array are available at an ESA server555https://euclid.esac.esa.int/msp/refdata/nisp/NISP-SPECTRO-PASSBANDS-V1. Available throughput numbers include contributions from the grisms, the telescope, and NISP Optical Assembly, as well as the mean detector quantum efficiency. For convenience, the tables are available in both ASCII and FITS format.
@@ Acknowledgements.
- Riva et al. (2011) Riva, M., Carty, J., Barriere, J., et al. 2011, in 14th European Space Mechanisms and Tribology Symposium - SPIE Vol. 8442 844211-13, 14
@@ Appendix A On ground calibration coefficients
Table 3 to 9 give the calibration coefficients for Eq. 3, used to model spatial dependency of the coefficient needed in Eq. 2.
| Grism | Coef. | Chebyshev polynomial coefficients | Regular polynomial coefficient | ||||||